A vector A→ has a magnitude of 44.0 m and points in a direction 26.0° above the positive x-axis. A second vector, B→, has a magnitude of 84.0 m and points in a direction 40.0° below the negative x-axis. Using the component method, find the magnitude of the vector →R=→A+→B
Answer #1

Ax = 44 cos 26 = 39.55 m

Ay = 44 sin 26 = 19.29 m

Bx = 84 cos 40 = 64.348 m

By = - 84 sin 40 = - 54 m

Horizontal component of resultant vector is given

Rx = 39.55 + 64.348 = 103.9 m

vertical component of resultant vector

Ry = 19.29 - 54 = - 34.71 m

magnitude

R^2 = rx^2 + Ry^2

R = 109.544 m

========

do comment in case any doubt will reply for sure.. Goodluck